# Dynamical Phase Transitions Across Slow and Fast Regimes in a Two-Tone Driven Duffing Resonator

### v1

@@ V.1 Dynamical phase transitions
Figure 4: The response amplitude $A(T)$ as a function of time for three modulation cycles. (a) Fixed low detuning $\Delta_{21}=0.42\Gamma$ and varying relative strength $h$, with the numbered yellow lines corresponding to $h=0.1$, $h=0.33$ and $h=0.65$ to mark the cuts shown in Fig. 3. (b) Same as (a) with fixed high detuning $\Delta_{21}=0.91\Gamma$. (c) Fixed low relative strength $h=0.36$ and varying detuning $\Delta_{21}$. (d) Same as (c) with fixed high relative strength $h=0.71$. Sharp dark green regions signal large amplitudes corresponding to jumps to the high amplitude state.
@@
How do the trajectories change when the modulation is not infinitely slow? To answer this, we numerically solve Eq. (12), initializing the system in the low-amplitude stationary state of the single-tone case ($h=0$) before introducing the secondary tone at $T=0$. We look at the dynamics at a fixed, finite detuning $\Delta_{21}$ to understand the effect of the relative strength $h$ on these transitions. The resulting amplitude dynamics $A(T)$ are contingent on $h$, see Fig. 3. For $h<h_{lb}$, the initial lower amplitude state is stable throughout the modulation and the system stays close to it, see Fig. 3(a). However, for $h>h_{lb}$, we see two distinct type of behaviours. In Fig. 3(b), the amplitude of the system rises above the low amplitude branch after it bifurcates, but never reaches the high-amplitude branch. In contrast, in Fig. 3(c), the system rises above the low-amplitude branch and reaches the high-amplitude branch during the modulation cycle.
@@
To look at the complex dependence of the dynamics on $\Delta_{21}$ and $h$, we fix one parameter at small and large values and sweep across the other, see Fig. 4. The sharp jumps in $A(T)$ for $h>h_{lb}$ in the slow beating limit have the same period as the drive modulation, see Fig. 4(a). However, as the beating becomes faster ($\Delta_{21}\sim\Gamma$) the response changes gradually and accumulates a phase lag, see Fig. 4(b). Furthermore, these dynamics are asymmetric with respect to $\Delta_{21}$: blue-detuned secondary tones ($\Delta_{21}>0$) trigger large amplitude variations over a wider range of detunings and lead to higher peak values than their red-detuned ($\Delta_{21}<0$) counterparts, see Fig. 4(c) and (d). The diverse nature of these amplitude dynamics across parameter space motivates using an effective quantity to distinguish the distinct dynamical regimes and track the transitions across these regimes.
@@
Figure 5: The average amplitude in one modulation cycle, $\bar{A}$ as a function of modulation depth $h$ and frequency $\Delta_{21}$. (a) Experimental data (cf. Appendix C)(b) Simulation using Eq. (12). (c) The simulation results for an extended detuning range $\Delta_{21}\in[-3\Gamma,3\Gamma]$ with the transition boundaries, black for bare resonance, blue for resonance corrected using low-amplitude state and pink for correction using high-amplitude state, plotted on top. The horizontal arrows correspond to the fixed $\Delta_{21}$ cuts and the vertical arrows correspond to fixed $h$ cuts in Fig. 4. The markers 11, 22 and 33 mark the three different behavioural regimes shown in Fig. 3.
@@ VI Phase Transition Boundaries
In Sec. V, we observed that the secondary tone can lead to modulation and instabilities of the stationary response to the primary tone. In the slow modulation limit, we used Eq. (12) to argue that the response at frequency $\Omega_{1}$ experiences a slow modulation due to the time-dependent effective drive amplitude in Eq. (13) and thus can exceed a threshold beyond which such single-tone response destabilizes. However, this approach does not take into account the time it takes the system to respond to changes in the amplitude of the external drive, of order $\Gamma^{-1}$. In Fourier domain, this entails that we neglected the fact that the resonator has a different susceptibility to the secondary tone compared to the primary tone, cf. Fig. 1(b). As such, we cannot account for why the threshold in Fig. 5 increases with $|\Delta_{21}|$, nor why this increase occurs in an asymmetric fashion with respect to $\Delta_{21}$.
@@
In this Section, we incorporate the information from linear response, i.e., the susceptibility function, on top of the stationary nonlinear motion of the Duffing resonator, to approximate the mechanism that allows for inter-attractor trajectories. We first propose that the two-tone linear response power absorbed by the resonator (6) serves as a good first approximation to the two-tone driven Duffing resonator. Secondly, we use this approximation to evaluate the filtered effective drive experienced at $\Omega_{1}$. To this end, we use Eq. (6) to calculate the maximum and minimum linear response power, $A^{2}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0},+$}}$ and $A^{2}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0},-$}}$ respectively, achieved in a modulation cycle. Using the identity, $\max_{t}[a_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0}$}}\cos(\Delta_{21}t)+b_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0}$}}\sin(\Delta_{21}t)]=\sqrt{a_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0}$}}^{2}+b_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0}$}}^{2}}$, the expressions for the extrema read
@@
No. Start branch Regime Outcome
1 lb ${F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{l},+$}}<F_{lb}$ stays lb
2 lb ${F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{l},+$}}<F_{lb}<{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{h},+$}}$ lb + deviation
3 lb $F_{lb}<{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{l},+$}},{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{h},+$}}\text{ and }{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{h},-$}}<F_{hb}$ lb $\leftrightarrow$ hb
4 lb $F_{lb}<{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{l},+$}},{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{h},+$}}\text{ and }{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{h},-$}}>F_{hb}$ moves to hb and stays
5 hb $F_{hb}<{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{l},-$}},{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{h},-$}}$ stays hb
6 hb ${F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{h},-$}}<F_{hb}<{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{l},-$}}$ hb + deviation
7 hb ${F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{l},-$}},{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{h},-$}}<F_{hb}\text{ and }{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{l},+$}},{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{h},+$}}>F_{lb}$ hb $\leftrightarrow$ lb
8 hb ${F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{l},-$}},{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{h},-$}}<F_{hb}\text{ and }{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{l},+$}},{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{h},+$}}<F_{lb}$ moves to lb and stays
Table 1: This table summarizes the different dynamical regions assuming a fixed primary tone in the bistability region, 1 - 8, differentiated by the starting branch [low(high) branch as lb(hb)] and the detuning and strength of the secondary tone. The properties of the secondary tone lead to different effective filtered drive strengths during modulation, see Eqs. (17) and (19), which we call regimes. $F_{hb}$, $F_{lb}$ are the bifurcation thresholds for the high and lower branch obtained in Eq. (11). ${F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{h},\pm$}}$ is the maximum (minimum) renormalized filtered effective drive obtained in Eq. 19 that takes into account the dispersive shift [cf. Eq. 18] due to the high amplitude state, $A_{\text{high}}$. ${F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{l},\pm$}}$ is the renormalized filtered effective drive that incorporates resonance renormalization due to the low amplitude state, $A_{\text{low}}$. 1 - 3 refer to the cases presented in Figs. 3(a)-(c) and marked on the phase diagram in Fig. 5(c).Cases 5- 7 are presented in Figs. 6(a)-(c) and are marked on the phase diagram in Fig. 7.
@@
The comparison between the effective drive strengths obtained in Eq. (17) and the critical drive amplitudes for bifurcation obtained in Eq. (11) provides us with the required force for jumping away from the lower branch towards the higher branch, i.e., ${F_{\mathrm{eff}}}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0},+$}}\geq F_{lb}$, cf. Fig. 2(c). The threshold condition ${F_{\mathrm{eff}}}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0},+$}}=F_{lb}$ leads to the black transition boundary in Fig. 5(c). In Sec. V, the slow modulation allowed the system to reach the upper branch simply by leaving the lower one. With fast modulation, however, we must check that the system can still absorb sufficient power from the secondary tone during the jump. Indeed, as the system leaves the lower branch and approaches the upper branch, the oscillation amplitude grows, and the resonance frequency shifts with amplitude according to [36]
$\Omega_{r}(A)=\Omega_{0}(1+\frac{3\alpha}{4\Omega_{0}^{2}}A^{2})\,$, (18)
@@
detuning the system from the secondary tone. This “dynamical” Duffing frequency shift (18) reduces the filtered effective drive strength (17) and potentially prevents the transition to the upper branch. Therefore, completing the transition requires enough power to leave the lower branch and overcome resonance renormalization while keeping the absorbed power, and thus the filtered force in Eq. (17) above the $F_{lb}$ threshold. We incorporate a conservative estimate of this effect using the amplitude of the upper branch, $A_{\text{high}}$, which is the target state of the transition, cf. Fig. 2(c), to calculate a renormalized filtered effective drive,
${F^{2}_{\mathrm{eff}}}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{h},\pm$}}=A^{2}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{h},\pm$}}/|\tilde{\chi}_{1}|^{2}\,$, (19)
@@
by combining Eqs. (16) and (18), defining the renormalized susceptibility $\tilde{\chi}_{1}=\chi_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{h}$}}(\Omega_{1})$ and the renormalized absorbed power extrema $A^{2}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{h},\pm$}}$ where $\Omega_{h}\equiv\Omega_{r}(A_{\mathrm{high}})$. The condition ${F_{\mathrm{eff}}}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{h},+$}}=F_{lb}$ yields the transition boundary for the system to leave the lower branch and reach the higher branch during the modulation cycle, see pink curve in Fig. 5(c).
@@
In summary, two distinct thresholds govern the transition dynamics: the black line in Fig. 5(c) identifies the condition where the lower branch becomes unstable and the system is driven away from it, while the pink line identifies the regime where absorbed power from the secondary tone remains sufficient to cross the lower-branch instability and reach the higher-branch, despite the amplitude-dependent frequency renormalization [cf. Eqs. (18) and (19)]. The interplay between these two thresholds explains the different dynamical regimes, which we illustrate with three representative trajectories: 1 lies below both thresholds, where the trajectory remains confined to the lower branch as it remains stable throughout the modulation cycle, see Fig. 3(a); 2 lies above the black but below the pink threshold, where the system departs from the lower branch but fails to reach the higher branch due to resonance frequency renormalization, see Fig. 3(b); and 3 lies above both thresholds, where the system possesses enough power to both destabilize the lower branch and complete the jump, successfully reaching the high-amplitude branch during its cycle, see Fig. 3(c). Importantly, the pink line provides a good approximation to the parabolic instability boundary observed in Fig. 5(c) for positive detuning ($\Delta_{21}>0$), marking the success of our analytical approach.
@@
Although our threshold estimates capture the main features of the two-tone dynamics, notable deviations arise for negative detuning ($\Delta_{21}<0$), where the black boundary departs significantly from the experimental/simulated transition. To improve this, we include resonance renormalization arising from the finite amplitude of the lower branch, $A_{\text{low}}$. We use Eq. (18) to define $\Omega_{l}\equiv\Omega_{r}(A_{\mathrm{low}})$ and correct Eq. (17) to get ${F_{\mathrm{eff}}}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{l},\pm$}}$. The resulting blue threshold line in Fig. 5(c) agrees more closely with the observed transition, especially at negative detunings. This points to the relevance of higher-order renormalization schemes, which could eventually reconcile the effective-drive model with the full nonlinear dynamics. Remaining discrepancies can be traced to (i) the coarse-grained nature of the time-averaged order parameter (15), which tends to smear transitions, and (ii) the simplifying assumptions underlying our analytical treatment, which are intended to capture the qualitative threshold rather than its exact position. The latter motivates the study of the problem using a multifrequency ansatz.

### v2

@@ V.1 Dynamical phase transitions
Figure 4: Time-domain response amplitude $A(T)$ over three modulation cycles. (a) Slow modulation regime ($\Delta_{21}=0.42\Gamma$): Numbered lines indicate cuts for $h=0.1$ (stable low), $h=0.33$ (deviation), and $h=0.65$ (switching), corresponding to the trajectories in Fig. 3. (b) Same as (a) with fixed high detuning $\Delta_{21}=0.91\Gamma$. The response exhibits significant phase lag and smooth changes compared with (a). (c) Fixed $h=0.36$ with varying detuning: $\Delta_{21}>0$ triggers larger amplitudes in response to the modulation than $\Delta_{21}<0$. (d) Same as (c) with $h=0.71$. Sharp dark green regions signal large amplitudes corresponding to jumps to the high amplitude state.
@@
How do the trajectories change when the modulation is not infinitely slow? To answer this, we numerically solve Eq. (12), initializing the system in the low-amplitude stationary state of the single-tone case ($h=0$) before introducing the secondary tone at $T=0$. We look at the dynamics at a fixed, finite detuning $\Delta_{21}$ to understand the effect of the relative strength $h$ on these transitions. The resulting amplitude dynamics $A(T)$ are contingent on $h$, see Fig. 3. For $h<h_{\mathrm{lb}}$, the initial lower amplitude state is stable throughout the modulation and the system stays close to it, see Fig. 3(a). However, for $h>h_{\mathrm{lb}}$, we see two distinct types of behaviours. In Fig. 3(b), the amplitude of the system rises above the low amplitude branch after it bifurcates, but never reaches the high-amplitude branch. In contrast, in Fig. 3(c), the system rises above the low-amplitude branch and reaches the high-amplitude branch during the modulation cycle.
@@
To look at the complex dependence of the dynamics on $\Delta_{21}$ and $h$, we fix one parameter at small and large values and sweep across the other, see Fig. 4. The sharp jumps in $A(T)$ for $h>h_{\mathrm{lb}}$ in the slow beating limit have the same period as the drive modulation, see Fig. 4(a). However, as the beating becomes faster ($\Delta_{21}\sim\Gamma$) the response accumulates a significant phase lag and changes smoothly, see Fig. 4(b). Furthermore, these dynamics are asymmetric with respect to $\Delta_{21}$: positively detuned secondary tones ($\Delta_{21}>0$) trigger large amplitude variations over a wider range of detunings and lead to higher peak values than their negatively detuned ($\Delta_{21}<0$) counterparts, see Fig. 4(c) and (d). The diverse nature of these amplitude dynamics across parameter space motivates using an effective quantity to distinguish the distinct dynamical regimes and track the transitions across these regimes.
@@
Figure 5: The average amplitude in one modulation cycle, $\bar{A}$, as a function of relative strength $h$ and frequency $\Delta_{21}$. (a) Experimental data (adapted from Ref. [31], cf. Appendix C). (b) Numerical simulation of Eq. (12). (c) Extended simulation range with analytical transition boundaries overlaid. The black line represents the instability threshold of the low-amplitude branch based on the bare resonance [cf. Eq. (17)]. The pink line represents the transition threshold incorporating the resonance frequency shift due to the high amplitude state [cf. Eq. (19)]. The blue line uses the renormalized frequency of the low-amplitude state [cf. Eq. (20)]. Markers 1- 3 correspond to the regimes illustrated in Fig. 3. The horizontal arrows correspond to the fixed $\Delta_{21}$ cuts and the vertical arrows correspond to fixed $h$ cuts in Fig. 4.
@@ VI Phase Transition Boundaries
In Sec. V, we observed that the secondary tone can lead to modulation and instabilities of the stationary response to the primary tone. In the slow modulation limit, we used Eq. (12) to argue that the response at frequency $\Omega_{1}$ experiences a slow modulation due to the time-dependent effective drive amplitude in Eq. (13) and thus can exceed a threshold beyond which the initial single-tone branch destabilizes. However, this approach does not take into account the time it takes the system to respond to changes in the amplitude of the external drive, of order $\Gamma^{-1}$. In Fourier domain, this approach entails that we neglected the fact that the resonator has a different susceptibility to the secondary tone compared to the primary tone, cf. Fig. 1(b). As such, we cannot account for why the threshold in Fig. 5 increases with $|\Delta_{21}|$, nor why this increase occurs in an asymmetric fashion with respect to $\Delta_{21}$.
No. Start branch Regime Outcome
1 lb ${F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{\mathrm{l}},+$}}<F_{\mathrm{lb}}$ stays lb
2 lb ${F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{\mathrm{l}},+$}}<F_{\mathrm{lb}}<{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{\mathrm{h}},+$}}$ lb + deviation
3 lb $F_{\mathrm{lb}}<{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{\mathrm{l}},+$}},{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{\mathrm{h}},+$}}\text{ and }{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{\mathrm{h}},-$}}<F_{\mathrm{hb}}$ lb $\leftrightarrow$ hb
4 lb $F_{\mathrm{lb}}<{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{\mathrm{l}},+$}},{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{\mathrm{h}},+$}}\text{ and }{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{\mathrm{h}},-$}}>F_{\mathrm{hb}}$ moves to hb and stays
5 hb $F_{\mathrm{hb}}<{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{\mathrm{l}},-$}},{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{\mathrm{h}},-$}}$ stays hb
6 hb ${F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{\mathrm{h}},-$}}<F_{\mathrm{hb}}<{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{\mathrm{l}},-$}}$ hb + deviation
7 hb ${F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{\mathrm{l}},-$}},{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{\mathrm{h}},-$}}<F_{\mathrm{hb}}\text{ and }{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{\mathrm{l}},+$}},{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{\mathrm{h}},+$}}>F_{\mathrm{lb}}$ hb $\leftrightarrow$ lb
8 hb ${F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{\mathrm{l}},-$}},{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{\mathrm{h}},-$}}<F_{\mathrm{hb}}\text{ and }{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{\mathrm{l}},+$}},{F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{\mathrm{h}},+$}}<F_{\mathrm{lb}}$ moves to lb and stays
Table 1: This table summarizes the different dynamical regions assuming a fixed primary tone in the bistability region, 1 - 8, differentiated by the starting branch [low(high) branch as lb(hb)] and the detuning and strength of the secondary tone. The properties of the secondary tone lead to different effective filtered drive strengths during modulation, see Eqs. (17) and (19), which we call regimes. $F_{\mathrm{hb}}$, $F_{\mathrm{lb}}$ are the bifurcation thresholds for the high and lower branch obtained in Eq. (11). ${F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{\mathrm{h}},\pm$}}$ is the maximum (minimum) renormalized filtered effective drive obtained in Eq. (19) that takes into account the dispersive shift [cf. Eq. (18)] due to the high amplitude state, $A_{\text{high}}$. ${F_{\mathrm{eff}}}_{\raisebox{-0.81375pt}{$\scriptscriptstyle\Omega_{\mathrm{l}},\pm$}}$ is the renormalized filtered effective drive that incorporates resonance renormalization due to the low amplitude state, $A_{\text{low}}$. 1 - 3 refer to the cases presented in Figs. 3(a)-(c) and marked on the phase diagram in Fig. 5(c).Cases 5- 7 are presented in Figs. 9(a)-(c) and are marked on the phase diagram in Fig. 10.
@@
In this Section, we incorporate the information from linear response, i.e., the susceptibility function, on top of the stationary nonlinear motion of the Duffing resonator, to approximate the mechanism that allows for inter-attractor trajectories. We first propose that the two-tone linear response power absorbed by the resonator, see Eq. (6) serves as a good first approximation to the two-tone driven Duffing resonator. Secondly, we use this approximation to evaluate the filtered effective drive experienced at $\Omega_{1}$. To this end, we use Eq. (6) to calculate the maximum and minimum linear response power, $A^{2}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0},+$}}$ and $A^{2}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0},-$}}$ respectively, achieved in a modulation cycle. Using the identity $\max_{T}[a_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0}$}}\cos(\Delta_{21}T)+b_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0}$}}\sin(\Delta_{21}T)]=\sqrt{a_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0}$}}^{2}+b_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0}$}}^{2}}$ at $T=\arctan(b_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0}$}}/a_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0}$}})/\Delta_{21}$, the expressions for the extrema read
@@
The comparison between the effective drive strengths obtained in Eq. (17) and the critical drive amplitudes for bifurcation obtained in Eq. (11) provides us with the required force for jumping away from the lower branch towards the higher branch, i.e., ${F_{\mathrm{eff}}}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0},+$}}\geq F_{\mathrm{lb}}$, cf. Fig. 2(c). The threshold condition ${F_{\mathrm{eff}}}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0},+$}}=F_{\mathrm{lb}}$ leads to the black transition threshold in Fig. 5(c). In Sec. V, the slow modulation allowed the system to reach the higher branch simply by leaving the lower one. With fast modulation, however, we must check that the system can still absorb sufficient power from the secondary tone during the jump. Indeed, as the system leaves the lower branch and approaches the higher branch, the oscillation amplitude grows, and the resonance frequency shifts with amplitude according to [36]
$\Omega_{\mathrm{r}}(A)=\Omega_{0}(1+\frac{3\alpha}{8\Omega_{0}^{2}}A^{2})\,$, (18)
@@
detuning the system from the secondary tone. This “dynamical” Duffing frequency shift (18) changes the filtered effective drive strength (17) and potentially prevents the transition to the higher branch. Therefore, completing the transition requires enough power to leave the lower branch and overcome resonance renormalization while keeping the absorbed power, and thus the filtered force in Eq. (17), above the $F_{\mathrm{lb}}$ threshold. We incorporate a conservative estimate of this effect using the amplitude of the higher branch, $A_{\text{high}}$, which is the target state of the transition once the lower branch destabilizes, cf. Fig. 2(c), to calculate a renormalized filtered effective drive,
${F^{2}_{\mathrm{eff}}}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{\mathrm{h}},\pm$}}=A^{2}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{\mathrm{h}},\pm$}}/|\chi_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{\mathrm{h}}$}}(\Omega_{1})|^{2}\,$, (19)
@@
by combining Eqs. (16) and (18), defining the renormalized susceptibility $\chi_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{\mathrm{h}}$}}(\Omega_{1})$ and the renormalized absorbed power extrema $A^{2}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{\mathrm{h}},\pm$}}$ where $\Omega_{\mathrm{h}}\equiv\Omega_{\mathrm{\mathrm{r}}}(A_{\mathrm{high}})$. The condition ${F_{\mathrm{eff}}}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{\mathrm{h}},+$}}=F_{\mathrm{lb}}$ yields the transition boundary for the system to leave the lower branch and reach the higher branch during the modulation cycle, see pink curve in Fig. 5(c).
@@
In summary, two distinct thresholds govern the transition dynamics: the black line in Fig. 5(c) identifies the condition where the lower branch becomes unstable and the system is driven away from it, while the pink line identifies the regime where absorbed power from the secondary tone remains sufficient to cross the lower-branch instability and reach the higher-branch, despite the amplitude-dependent frequency renormalization [cf. Eqs. (18) and (19)]. The interplay between these two thresholds explains the different dynamical regimes, which we illustrate with three representative trajectories: 1 lies below both thresholds, where the trajectory remains confined to the lower branch as it remains stable throughout the modulation cycle, see Fig. 3(a); 2 lies above the black but below the pink threshold, where the system departs from the lower branch but fails to reach the higher branch due to resonance frequency renormalization, see Fig. 3(b); and 3 lies above both thresholds, where the system possesses enough power to both destabilize the lower branch and complete the jump, successfully reaching the higher branch during its cycle, see Fig. 3(c). Importantly, the pink line provides a good approximation to the parabolic instability boundary observed in Fig. 5(c) for positive detuning ($\Delta_{21}>0$), marking the success of our analytical approach.
Although our threshold estimates capture the main features of the two-tone dynamics, notable deviations arise for negative detuning ($\Delta_{21}<0$), where the black transition threshold deviates significantly from the experimental/simulated transition. To improve this, we include resonance renormalization arising from the finite amplitude of the lower branch, $A_{\text{low}}$ close to the system’s initial state. We use Eq. (18) to define $\Omega_{\mathrm{l}}\equiv\Omega_{\mathrm{r}}(A_{\mathrm{low}})$ and correct Eq. (17) to get
${F_{\mathrm{eff}}}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{\mathrm{l}},\pm$}}=A^{2}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{\mathrm{l}},\pm$}}/|\chi_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{\mathrm{l}}$}}(\Omega_{1})|^{2}\,$, (20)
@@
the renormalized filtered effective drive incorporating the shift due to $A_{\mathrm{low}}$. The resulting blue threshold line in Fig. 5(c) is closer to the observed transition for negative detunings ($\Delta_{21}<0$). This points to the relevance of higher-order renormalization schemes, which could eventually reconcile the effective-drive model with the full nonlinear dynamics. Remaining discrepancies can be traced to (i) the coarse-grained nature of the time-averaged order parameter (15), which tends to smear transitions, and (ii) the simplifying assumptions underlying our analytical treatment, which are intended to capture the qualitative threshold rather than its exact position. The latter motivates the study of the problem using a multifrequency ansatz.